# Learning Repetition-Invariant Representations for Polymer Informatics

### v1

@@ Abstract
Polymers are large macromolecules composed of repeating structural units known as monomers and are widely applied in fields such as energy storage, construction, medicine, and aerospace. However, existing graph neural network methods, though effective for small molecules, only model the single unit of polymers and fail to produce consistent vector representations for the true polymer structure with varying numbers of units. To address this challenge, we introduce Graph Repetition Invariance (GRIN), a novel method to learn polymer representations that are invariant to the number of repeating units in their graph representations. GRIN integrates a graph-based maximum spanning tree alignment with repeat-unit augmentation to ensure structural consistency. We provide theoretical guarantees for repetition-invariance from both model and data perspectives, demonstrating that three repeating units are the minimal augmentation required for optimal invariant representation learning. GRIN outperforms state-of-the-art baselines on both homopolymer and copolymer benchmarks, learning stable, repetition-invariant representations that generalize effectively to polymer chains of unseen sizes.
@@ 1 Introduction
Polymers are materials composed of macromolecules made up of multiple repeating units (RUs), i.e., monomers with polymerization points. For example, a homopolymer consists of a single RU repeated along the chain (e.g., polyethylene), resulting in uniform properties. A copolymer interleaves two or more distinct RUs (e.g., styrene-butadiene rubber) to combine or tune material characteristics [5]. Figure 1 illustrates two examples of polymers, which can be encoded as graphs of atoms and bonds. Polymer informatics is an emerging discipline at the intersection of materials science and machine learning, aiming to accelerate the discovery and design of new polymers through data-driven approaches [2, 6, 18, 22], such as graph representation learning. In this direction, graph neural networks (GNNs) have been developed for small-molecule tasks [4, 13, 31, 33]; however, their generalization to polymers has been limited to a single repeating unit.
@@
Figure 2 (a) visualizes the t-SNE embeddings generated by a standard two-layer GCN, trained with and without augmentation, evaluated on test sets augmented with up to 20 RUs. For the same polymer, embeddings, where darker colors indicate larger repeat sizes in its graphs, drift across the latent space as the number of repeats increases. This suggests that conventional message passing entangles repeat size with the core chemical structure of the repeating unit, failing to produce repetition-invariant representations. GCN benefits from repeat-unit augmentation during training, as embeddings of the same polymer with smaller repeat sizes begin to coalesce into visible striations. However, the clusters remain dispersed overall, and embeddings for larger repeat sizes are still scattered, indicating that augmentation alone is not efficient in achieving true generalization.
@@
In this work, we propose Graph Repetition INvariance (GRIN), a method that combines algorithm alignment with repeat-unit augmentation to address the challenge of repetition-invariant representation learning. GRIN aligns message passing with the edge-greedy logic of the maximum spanning tree (MST) algorithm: dynamic programming(DP)-style updates enforce a fixed reasoning procedure, enabling intrinsic generalization across graph sizes. The MST criterion further preserves the most informative backbone of the polymer graph. We provide theory guarantees for invariant representation learning both model-wise and data-wise: Model-wise, GRIN’s max-aggregation with a sparsity penalty aligns it with the MST construction process, inheriting the size-generalization guarantees from DP. Data-wise, under the repeating unit contraction (definition 3.1), we formally establish latent repetition-invariance (proposition 3.2) and identify the minimal repeat size required for optimal invariant learning (proposition 3.3).
@@ 2.1 Invariant Representation Learning
Current invariant graph representation learning focuses on identifying features that remain stable across different environments to achieve out-of-distribution (OOD) generalization on graphs [8, 16, 19]. Early work extended invariant risk minimization to graphs, seeking features that remain predictive across training environments [3]. DIR [32] masks environment-specific subgraphs via counterfactual interventions, while Liu et al. [18] proposed a method GREA to learn graph rationales with environment-based augmentations by swapping the environment subgraphs between different samples. Although effective against covariate and local-structure shifts, these approaches largely ignore distributional changes in graph scale.
Size generalization, a specific instantiation of OOD generalization, aiming at learning consistent representations when node or edge counts vary. To this end, SizeShiftReg (SSR) [7] enforces invariance by coarsening graphs during training and aligning embeddings of original and reduced-scale counterparts. Huang et al. [14] proposed a disentanglement loss that separates size-related factors from task-specific features. However, our method explicitly targets the unique challenge of polymers, where causal information must be preserved as graph size changes due to varying numbers of repeating units.
@@ 2.2 Neural Algorithmic Alignment
Neural algorithmic alignment has recently emerged as a powerful tool for enhancing GNN size extrapolation [9, 11, 28]. In classical algorithms like Bellman-Ford [10] and Prim [23], solutions are typically constructed through a series of iterations. Neural algorithmic alignment enforces each message-passing layer to mirror one iteration of a known graph algorithm, rather than treating the network as a black-box end-to-end mapper. Although this approach yields impressive OOD performance on well-defined algorithmic benchmarks, its scope remains limited. For instance, Nerem et al. [21] demonstrated provable extrapolation for shortest-path problem by aligning a GNN with Bellman–Ford. Existing models remain restricted to synthetic benchmarks [29] and do not tackle real-world graph classification or regression tasks.
@@ 3 Methodology
In this section we present Graph Repetition Invariance (GRIN), a novel framework that i) augments polymer graphs by chaining repeating units and ii) aligns message passing with a maximum-spanning-tree (MST) on that chain. GRIN learns representations invariant to the number of repeating units.
@@ 3.1 Problem Definition
All graphs in the family , share the same RU and an identical ground-truth property value. Training on enforces a learning objective in which the model must predict the same property regardless of the repeat count.
We propose a repeat-unit augmentation strategy that generates multiple graph instances of a polymer with varying numbers of RUs. Each RU represents either a homopolymer monomer (-A-) or a copolymer motif (-A-B-). Real polymers often contain a very large number (-) of such RUs, far beyond the scope of traditional methods can learn. It is also infeasible to directly augment the training set with all possible repeat sizes. Therefore, in Section 3.3.2, we analyze the minimal repeat size required to learn invariant representations for this augmentation strategy.
@@ 3.2 Message Passing with Max Aggregation
and are multi-layer perceptrons (MLPs) and is element-wise maximum over incoming messages. With residual connection , any node not chosen by max-aggregation simply skips the update and retains its previous embedding unchanged. The function takes the features of neighbor and the edge into a message vector. updates the embedding of node . Max-aggregation enforces a sparsity bias in the aggregation step, allowing only the most informative neighbor to influence each node update that mirrors the edge-pruning principle of MST.
@@ 3.3 Theoretical Guarantee of Invariant Representation Learning
In this section, we provide GRIN’s theoretical guarantees for repetition-invariant learning. First, we show that with max aggregation and sparsity constraint, GRIN emulates a greedy DP recurrence and inherits its size-generalization properties. Second, we prove that augmenting with three RUs is sufficient to reach optimal repetition invariance.
@@ 3.3.1 Model-Wise: Algorithm Alignment with Maximum-Spanning-Tree
where represents the edge weight, given our message passing design in Eq. 1, we have
The additive residual term in Eq. 1 ensures that only the node selected by max-aggregation at layer receives an update, while all other nodes satisfy , perfectly mirroring Prim’s algorithm where nodes not chosen in the greedy step remain unchanged. Additionally, we incorporate an sparsity penalty on model parameters as
encourages many entries of to become exactly zero. This further reinforces the MST alignment by pruning weak connections in the learned message scores, so that only the strongest (MST-relevant) edges survive, enhancing both sparsity and interpretability.
@@
While max-aggregation selects the single most informative incoming message at each node, classical sum- and mean-aggregation mix all neighbor messages[17]. In particular, for sum and mean we have:
Sum-aggregation grows linearly with the node degree, and mean-aggregation normalizes by degree but still blends every neighbor’s contribution. Both operations depend on the size of the neighborhood:
In contrast, by combining max-aggregation with an sparsity penalty, GRIN effectively emulates MST construction and thus inherits its inherent size-generalization capabilities, as aligning to DP recurrences enjoy out-of-distribution generalization across input scales [35].
@@ 3.3.2 Data-Wise: Minimal Repeat Size for Invariant Representation
One can repeat the polymer unit (-A-) as the augmented data point , it can be viewed as an undirected graph . Let represents the number of repeats (or the degree of polymerization). We analyze the model behavior over an abstract hyperchain level.
@@
We contract each repeating unit -A- in the polymer graph to a single supernode and neglecting all intra-monomer edges, while retaining only the inter-monomer polymerization edges. can be abstracted as a hyperchain
@@
We denote by the embedding of supernode after the -th layer, and as its aggregated message. Then layer-wise message passing reduces to
Given Eq. 5, we notice that training only on (hyperchain of length 1) activates solely the branch, leaving parameters unlearned which is essential in extrapolating to longer chains. To address this issue, we supervise on a merge set with a shared target . Let loss be written as:
Here is supernode ’s embedding after layers, and is the term introduced in Eq. 3.
Proposition 3.2 (Latent Repetition-Invariance).
Under definition 3.1 and let , for every test hyperchain with , the prediction is
Training on exposes both structural cases—nodes of and —and the term zeros out all other pathways, which means the complete update rule in Eq. 5 is learned. Consequently, in () every node has , the model applies the same update rule at each layer, yielding a constant output , irrespective of the chain length .
@@
Introducing augmentation with supernodes satisfying helps model learn extrapolation across repeating units. A hyperdegree-1 supernode receives exactly one incoming message, so its gradient flows along a single branch. In contrast, a hyperdegree-2 supernode has two competing branches, introducing supervision for multi-branch aggregation.
Let be the back-propagated error arriving at a hyperdegree-2 supernode with two neighbors . Under classical GNN contraction assumption[26], each message–passing layer is -Lipschitz with , so any gradient norm is reduced by at most a factor of per layer.
@@
When supervising on with , the total back-propagated gradient norm at any degree-2 supernode satisfies:
Under the hyperchain abstraction, (the smallest sample containing a hyperdegree-2 node) is the minimal configuration for GRIN to learn true multi-branch fusion. never exposes this dual-message scenario and thus could be insufficient. Furthermore, given the contraction factor , increasing repeat size beyond 3 yields only geometrically vanishing improvements with additional training cost.
@@ 4 Experiments
We conduct experiments for three research questions: Q1) Effectiveness: Does GRIN make more accurate predictions for homopolymers and copolymers than existing methods? Q2) Repetition-Invariance Learning: Does GRIN learn invariant representations and maintain performance across varying repeat sizes? Q3) Ablation study: What is the minimal repeat size required for augmentation? How does the merge ratio (proportion of augmented graphs) impact performance?
@@ 4.1 Experimental Settings
We evaluate on four homopolymer datasets and two copolymer datasets. The four datasets predict the glass transition temperature (GlassTemp, ∘C), polymer density (PolyDensity, g/cm3), melting temperature (MeltingTemp, ∘C) and oxygen permeability (O2Perm, Barrer). The other two are about electron affinity (EA, eV) and ionization potential (IP, eV). The dataset statistics are given in Table 1, in which the Size refers to the diameter of graph calculated by
@@
For every polymer, we construct a family of polymer graphs as Section 3.1. We use GRIN-RepAug to denote model training on and GRIN for model training with repeat-unit augmentation ,. Dataset details can be found in the appendix A.
@@
We evaluate the regression performance using the coefficient of determination (R2) and Root Mean Squared Error (RMSE). For baselines, we compare our GRIN with methods designed specifically for size-generalization: DISGEN, SSR, BFGNN [7, 14, 21] and others for general OOD, including GREA, DIR, IRM, RPGNN [3, 18, 20, 32]. We test both GIN and GCN as graph encoder for all models. Please refer to appendix B for details of implementation.
@@ 4.2 Results on Effectiveness (Q1) and Repetition-Invariant Representation (Q2)
We evaluate on for homopolymer and for copolymer, graph diameter ranges from 10 to 104. GRIN consistently achieves best results among all tasks, GRIN-RepAug ranks the second, which demonstrate the method’s effectiveness of repetition-invariant learning. All observations hold for both GCN and GIN backbones, confirming that the gain stems from the repeat-aware alignment itself rather than the choice of encoder. More results can be found at Section C.1.
@@
Table 2 and Table 3 report R2 and RMSE on Test1 (1 RU) and Test60 (60 RUs). Test1 measures property prediction performance, whereas Test60 probes a 60 repeat-size extrapolation. Effectiveness: Across all four homopolymer tasks, GRIN with/without augmentation achieves top-2 accuracy on all test sets. Compared to the strongest baseline, GRIN achieves a 1-2% improvement on GlassTemp, MeltingTemp, and O2Perm, and a remarkable 15% on PolyDensity. The advantage widens under repeat-size extrapolation: the improvement exceeds 10% for Test60. Several baselines collapse on Test60 (negative R2). In particular, DISGEN almost fails on PolyDensity and O2Perm: its augmentation pipeline depends on a GNN-Explainer module that can misprioritize nodes (common in the noisy, low-sample O2Perm and PolyDensity datasets), results in a training failure. Repetition Invariance: GRIN’s performance gap between Test1 and Test60 stays within ±3%, confirming that the model indeed learns repetition-invariant representation. Once trained on 1-RU and 3-RU graphs, it transfers to polymers sixty times larger without loss of accuracy.
@@
Table 4 summarizes results on two copolymer datasets. Effectiveness: GRIN ranks first in every setting and GRIN-RepAug the second. Against the strongest baseline, GRIN raises R2 by 1.8-3.7% and cuts RMSE by 14-30% on Test1. The margin widens on Test20, reaching 5.2-7.4% (R2) and 24-35% (RMSE). Baselines show the same sharp performance decline on larger graphs as observed in the homopolymer experiments, underscoring the challenge of repeat-size extrapolation. Repetition Invariance: When graph size grows 20, GRIN’s accuracy remains nearly unchanged: R2 drops by at most 0.5% and RMSE increases by less than 2%.
Model EA IP Test1 Test20 Test1 Test20 R2 RMSE R2 RMSE R2 RMSE R2 RMSE GCN GCN [15] 0.939±0.006 0.146±0.008 0.887±0.023 0.199±0.020 0.918±0.004 0.141±0.004 0.909±0.004 0.149±0.003 DIR [32] < 0 > 1 < 0 > 1 < 0 > 1 < 0 > 1 GREA [18] 0.939±0.007 0.146±0.008 < 0 > 1 0.905±0.069 0.146±0.053 < 0 > 1 IRM [3] 0.930±0.012 0.157±0.013 < 0 > 1 < 0 > 1 RPGNN [20] 0.929±0.003 0.158±0.004 < 0 > 1 0.914±0.020 0.144±0.017 < 0 > 1 SSR [7] 0.818±0.091 0.248±0.062 < 0 > 1 0.883±0.021 0.168±0.015 < 0 > 1 DISGEN [14] 0.163±0.081 0.542±0.027 0.179±0.082 0.537±0.027 0.232±0.047 0.432±0.013 0.174±0.062 0.447±0.017 BFGCN [21] 0.904±0.014 0.184±0.014 0.883±0.018 0.203±0.016 0.916±0.014 0.137±0.006 0.901±0.014 0.146±0.008 GRIN-RepAug 0.952±0.005 0.129±0.007 0.942±0.006 0.143±0.008 0.946±0.010 0.113±0.013 0.936±0.010 0.125±0.010 GRIN 0.956±0.002 0.125±0.002 0.952±0.001 0.129±0.002 0.952±0.005 0.108±0.006 0.947±0.005 0.113±0.005 GIN GIN [34] 0.918±0.006 0.170±0.006 0.903±0.007 0.185±0.007 0.925±0.004 0.135±0.004 0.914±0.004 0.144±0.003 DIR [32] < 0 > 1 < 0 > 1 < 0 > 1 < 0 > 1 GREA [18] 0.831±0.091 0.237±0.072 < 0 > 1 0.918±0.002 0.141±0.002 < 0 > 1 IRM [3] 0.930±0.010 0.157±0.011 < 0 > 1 0.934±0.003 0.126±0.003 < 0 > 1 RPGNN [20] 0.929±0.001 0.158±0.002 < 0 > 1 0.917±0.025 0.141±0.021 < 0 > 1 SSR [7] 0.810±0.161 0.245±0.104 < 0 > 1 0.860±0.067 0.182±0.042 < 0 > 1 DISGEN [14] 0.123±0.146 0.554±0.048 0.100±0.107 0.565±0.035 0.232±0.047 0.432±0.013 0.181±0.060 0.447±0.016 BFGIN [21] 0.925±0.006 0.163±0.007 0.918±0.004 0.170±0.004 0.929±0.006 0.131±0.006 0.918±0.002 0.141±0.002 GRIN-RepAug 0.954±0.004 0.127±0.006 0.947±0.004 0.137±0.005 0.963±0.001 0.095±0.001 0.953±0.003 0.107±0.004 GRIN 0.961±0.001 0.117±0.002 0.960±0.001 0.119±0.002 0.965±0.001 0.092±0.001 0.962±0.001 0.096±0.001

### v2

@@ Abstract
Polymers are large macromolecules composed of repeating structural units known as monomers and are widely applied in fields such as energy storage, construction, medicine, and aerospace. However, existing graph neural network methods, though effective for small molecules, only model the single unit of polymers and fail to produce consistent vector representations for the true polymer structure with varying numbers of units. To address this challenge, we introduce Graph Repetition Invariance (GRIN), a novel method to learn polymer representations that are invariant to the number of repeating units in their graph representations. GRIN integrates a graph-based maximum spanning tree alignment with repeat-unit augmentation to ensure structural consistency. We provide theoretical guarantees for repetitioninvariance from both model and data perspectives, demonstrating that three repeating units are the minimal augmentation required for optimal invariant representation learning. GRIN outperforms state-of-the-art baselines on both homopolymer and copolymer benchmarks, learning stable, repetition-invariant representations that generalize effectively to polymer chains of unseen sizes.
@@ 1 Introduction
Polymers are materials composed of macromolecules made up of multiple repeating units (RUs), i.e., monomers with polymerization points. For example, a homopolymer consists of a single RU repeated along the chain (e.g., polyethylene), resulting in uniform properties. A copolymer interleaves two or more distinct RUs (e.g., styrene‑butadiene rubber) to combine or tune material characteristics [5]. Figure˜1 illustrates two examples of polymers, which can be encoded as graphs of atoms and bonds. Polymer informatics is an emerging discipline at the intersection of materials science and machine learning, aiming to accelerate the discovery and design of new polymers through data-driven approaches [2, 6, 18, 23, 19], such as graph representation learning. In this direction, graph neural networks (GNNs) have been developed for small-molecule tasks [4, 13, 34, 36]; however, their generalization to polymers has been limited to a single repeating unit.
@@
Figure˜2 (a) visualizes the t-SNE embeddings generated by a standard two-layer GCN, trained with and without augmentation, evaluated on test sets augmented with up to 20 RUs. For the same polymer, embeddings, where darker colors indicate larger repeat sizes in its graphs, drift across the latent space as the number of repeats increases. This suggests that conventional message passing entangles repeat size with the core chemical structure of the repeating unit, failing to produce repetition-invariant representations. GCN benefits from repeat-unit augmentation during training, as embeddings of the same polymer with smaller repeat sizes begin to coalesce into visible striations. However, the clusters remain dispersed overall, and embeddings for larger repeat sizes are still scattered, indicating that augmentation alone is not efficient in achieving true generalization.
@@
In this work, we propose Graph Repetition INvariance (GRIN), a method that combines algorithm alignment with repeat-unit augmentation to address the challenge of repetition-invariant representation learning. GRIN draws inspiration from the edge-greedy logic from the maximum spanning tree (MST) algorithm: dynamic programming(DP)-style updates enforce a fixed reasoning procedure, enabling intrinsic generalization across graph sizes. The MST criterion further preserves the most informative backbone of the polymer graph. We provide theory guarantees for invariant representation learning both model-wise and data-wise: Model-wise, GRIN’s max-aggregation with a sparsity penalty heuristically encourages alignment with the MST construction process, inheriting the size‑generalization guarantees from DP. Data-wise, under the repeating unit contraction (definition˜3.1), we formally establish latent repetition‑invariance (proposition˜3.2) and identify the minimal repeat size required for optimal invariant learning (proposition˜3.3).
@@ 2.1 Invariant Representation Learning
Current invariant graph representation learning focuses on identifying features that remain stable across different environments to achieve out‑of‑distribution (OOD) generalization on graphs [8, 16, 20]. Early work extended invariant risk minimization to graphs, seeking features that remain predictive across training environments [3]. DIR [35] masks environment-specific subgraphs via counterfactual interventions, while Liu et al. [18] proposed a method GREA to learn graph rationales with environment-based augmentations by swapping the environment subgraphs between different samples. Although effective against covariate and local‑structure shifts, these approaches largely ignore distributional changes in graph scale.
Size generalization, a specific instantiation of OOD generalization, aiming at learning consistent representations when node or edge counts vary. To this end, SizeShiftReg (SSR) [7] enforces invariance by coarsening graphs during training and aligning embeddings of original and reduced-scale counterparts. Huang et al. [14] proposed a disentanglement loss that separates size‑related factors from task‑specific features. However, our method explicitly targets the unique challenge of polymers, where key structural information must be preserved as graph size changes due to varying numbers of repeating units.
@@ 2.2 Neural Algorithmic Alignment
Neural algorithmic alignment has recently emerged as a powerful tool for enhancing GNN size extrapolation [9, 11, 30]. In classical algorithms like Bellman-Ford [10] and Prim [24], solutions are typically constructed through a series of iterations. Neural algorithmic alignment enforces each message‑passing layer to mirror one iteration of a known graph algorithm, rather than treating the network as a black‑box end‑to‑end mapper. Although this approach yields impressive OOD performance on well‑defined algorithmic benchmarks, its scope remains limited. For instance, Nerem et al. [22] demonstrated provable extrapolation for shortest‑path problem by aligning a GNN with Bellman–Ford. Existing models remain restricted to synthetic benchmarks [31] and do not tackle real-world graph classification or regression tasks.
@@ 3 Methodology
In this section we present Graph Repetition Invariance (GRIN), a novel framework that i) augments polymer graphs by chaining repeating units and ii) aligns message passing with a maximum‐spanning‐tree (MST) on that chain. GRIN learns representations invariant to the number of repeating units.
@@ 3.1 Problem Definition
All graphs in the family , share the same RU and an identical ground‑truth property value. Training on enforces a learning objective in which the model must predict the same property regardless of the repeat count.
We propose a repeat‐unit augmentation strategy that generates multiple graph instances of a polymer with varying numbers of RUs. Each RU represents either a homopolymer monomer (-A-) or a copolymer motif (-A-B-). Real polymers often contain a very large number (-) of such RUs, far beyond the scope of traditional methods can learn. It is also infeasible to directly augment the training set with all possible repeat sizes. Therefore, in Section˜3.3.2, we analyze the minimal repeat size required to learn invariant representations for this augmentation strategy.
@@ 3.2 Message Passing with Max Aggregation
and are multi-layer perceptrons (MLPs) and is element-wise maximum over incoming messages. With previous state , any non‑max neighbour contributes zero to leaves corresponding feature dimensions remain unchanged. The function takes the features of neighbor and the edge into a message vector. updates the embedding of node . Max aggregation imposes a selection bias toward the strongest neighbor, we add an L1 sparsity penalty on the message and update networks to suppress non‑dominant pathways (Eq. 3), encouraging a sparse, MST‑like backbone.
@@ 3.3 Theoretical Guarantee of Invariant Representation Learning
In this section, we provide GRIN’s theoretical guarantees for repetition-invariant learning. First, we show that with max aggregation and sparsity constraint, GRIN emulates a greedy DP recurrence and inherits its size‑generalization properties. Second, we prove that augmenting with three RUs is sufficient to reach optimal repetition invariance.
@@ 3.3.1 Model-Wise: Algorithm Alignment with Maximum-Spanning-Tree
where represents the edge weight, given our message passing design in Eq.˜1, we have
The additive residual term in Eq.˜1 ensures that only the feature selected by max‑aggregation at layer receives an update, while all othes satisfy , mirroring Prim’s algorithm where nodes not chosen in the greedy step remain unchanged. Additionally, we incorporate an sparsity penalty on model parameters as
encourages many entries of to become exactly zero. This further reinforces the MST alignment by pruning weak connections in the learned message scores, so that only the strongest (MST‑relevant) edges survive, enhancing both sparsity and interpretability.
@@
While max‑aggregation selects the single most informative incoming message at each node, classical sum‑ and mean‑aggregation mix all neighbor messages[17]. In particular, for sum and mean we have:
Sum‑aggregation grows linearly with the node degree, and mean‑aggregation normalizes by degree but still blends every neighbor’s contribution. Both operations depend on the size of the neighborhood:
These findings are further corroborated by the experimental results in Table˜12, which show that sum- and mean-aggregation degrade substantially in performance as the repeat size increases. In contrast, by combining max‑aggregation with an sparsity penalty, GRIN effectively emulates MST construction and thus inherits its inherent size‑generalization capabilities, as aligning to DP recurrences enjoy out‐of‐distribution generalization across input scales [38].
@@ 3.3.2 Data-Wise: Minimal Repeat Size for Invariant Representation
One can repeat the polymer unit (-A-) as the augmented data point , it can be viewed as an undirected graph . Let represents the number of repeats (or the degree of polymerization). We analyze the model behavior over an abstract hyperchain level and show that repeat-augmented polymers ( 2), the model can apply same update rule at each layer, yielding a constant output , irrespective of the repeat size.
@@
We contract each repeating unit -A- in the polymer graph to a single supernode and neglecting all intra‑monomer edges, while retaining only the inter‑monomer polymerization edges. can be abstracted as a hyperchain
@@
We denote by the embedding of supernode after the ‑th layer, and as its aggregated message. Then layer‐wise message passing reduces to
Given Eq.˜5, we notice that training only on (hyperchain of length 1) activates solely the branch, leaving parameters unlearned which is essential in extrapolating to longer chains. To address this issue, we supervise on a merge set with a shared target . Let loss be written as:
Here is supernode ’s embedding after layers, and is the term introduced in Eq.˜3.
Proposition 3.2 (Latent Repetition‐Invariance).
Under definition˜3.1 and let , for every test hyperchain with , the prediction is
Training on exposes both structural cases—nodes of and —and the term zeros out all other pathways, which means the complete update rule in Eq.˜5 is learned.
@@
Introducing augmentation with supernodes satisfying helps model learn extrapolation across repeating units. A hyperdegree‑1 supernode receives exactly one incoming message, so its gradient flows along a single branch. In contrast, a hyperdegree-2 supernode has two competing branches, introducing supervision for multi‑branch aggregation.
Let be the back‑propagated error arriving at a hyperdegree‑2 supernode with two neighbors . Under classical GNN contraction assumption[28], each message–passing layer is assumed to be -Lipschitz, i.e., it satisfies for all inputs , and some constant <1, so any gradient norm is reduced by at most a factor of per layer.
@@
When supervising on with , the total back‑propagated gradient norm at any degree‑2 supernode satisfies:
Under the hyperchain abstraction, (the smallest sample containing a hyperdegree‑2 node) is the minimal configuration for GRIN to learn true multi‑branch fusion. never exposes this dual‑message scenario and thus could be insufficient. Furthermore, given the contraction factor , increasing repeat size beyond 3 yields only geometrically vanishing improvements with additional training cost.
@@ 4 Experiments
We conduct experiments for three research questions:
Q1) Effectiveness: Does GRIN make more accurate predictions for homopolymers and copolymers than existing methods?
Q2) Repetition-Invariance Learning: Does GRIN learn invariant representations and maintain performance across varying repeat sizes?
Q3) Ablation study: What is the minimal repeat size required for augmentation? How does the merge ratio (proportion of augmented graphs) impact performance?
@@ 4.1 Experimental Settings
We evaluate property prediction task on four homopolymer datasets and two copolymer datasets. The four datasets predict the glass transition temperature (GlassTemp, ∘C), polymer density (PolyDensity, g/cm3), melting temperature (MeltingTemp, ∘C) and oxygen permeability (O2Perm, Barrer). The other two are about electron affinity (EA, eV) and ionization potential (IP, eV). The dataset statistics are given in Table˜1, in which the Size refers to the diameter of graph calculated by
@@
For every polymer, we construct a family of polymer graphs as Section˜3.1. We use GRIN-RepAug to denote model training on and GRIN for model training with repeat-unit augmentation ,. Dataset details can be found in the appendix˜A.
@@
We evaluate the regression performance using the coefficient of determination (R2) and Root Mean Squared Error (RMSE). For baselines, we compare our GRIN with methods designed specifically for size-generalization: DISGEN, SSR, BFGNN [7, 14, 22] and others for general OOD, including GREA, DIR, IRM, RPGNN [3, 18, 21, 35]. We test both GIN and GCN as graph encoder for all models. Please refer to appendix˜B for details of implementation and appendix˜C for computational efficiency comparison.
@@ 4.2 Results on Effectiveness (Q1) and Repetition-Invariant Representation (Q2)
We evaluate on for homopolymer and for copolymer, graph diameter ranges from 10 to 104. GRIN consistently achieves best results among all tasks, GRIN-RepAug ranks the second, which demonstrate the method’s effectiveness of repetition-invariant learning. All observations hold for both GCN and GIN backbones, confirming that the gain stems from the repeat-aware alignment itself rather than the choice of encoder. More results can be found at appendix˜D.
@@
Table 2 and Table 3 report R2 and RMSE on Test1 (1 RU) and Test60 (60 RUs). Test1 measures property prediction performance, whereas Test60 probes a 60 repeat-size extrapolation.
Effectiveness: Across all four homopolymer tasks, GRIN with/without augmentation achieves top-2 accuracy on all test sets. Compared to the strongest baseline, GRIN achieves a 1-2% improvement on GlassTemp, MeltingTemp, and O2Perm, and a remarkable 15% on PolyDensity. The advantage widens under repeat-size extrapolation: the improvement exceeds 10% for Test60. Several baselines collapse on Test60 (negative R2). In particular, DISGEN almost fails on PolyDensity and O2Perm: its augmentation pipeline depends on a GNN-Explainer module that can misprioritize nodes (common in the noisy, low-sample O2Perm and PolyDensity datasets), results in a training failure.
Repetition Invariance: GRIN’s performance gap between Test1 and Test60 stays within ±3%, confirming that the model indeed learns repetition-invariant representation. Once trained on 1-RU and 3-RU graphs, it transfers to polymers sixty times larger without loss of accuracy.
@@
Table 4 reports R2 and RMSE on Test1 (1 RU) and Test20 (20 RUs). Test1 measures property prediction performance, whereas Test20 probes a 20 repeat-size extrapolation.
Effectiveness: GRIN ranks first in every setting and GRIN-RepAug the second. Against the strongest baseline, GRIN raises R2 by 1.8-3.7% and cuts RMSE by 14-30% on Test1. The margin widens on Test20, reaching 5.2-7.4% (R2) and 24-35% (RMSE). Baselines show the same sharp performance decline on larger graphs as observed in the homopolymer experiments, underscoring the challenge of repeat‑size extrapolation.
Repetition Invariance: When graph size grows 20, GRIN’s accuracy remains nearly unchanged: R2 drops by at most 0.5% and RMSE increases by less than 2%.
Model EA IP Test1 Test20 Test1 Test20 R2 RMSE R2 RMSE R2 RMSE R2 RMSE GCN GCN [15] 0.939±0.006 0.146±0.008 0.887±0.023 0.199±0.020 0.918±0.004 0.141±0.004 0.909±0.004 0.149±0.003 DIR [35] < 0 > 1 < 0 > 1 < 0 > 1 < 0 > 1 GREA [18] 0.939±0.007 0.146±0.008 < 0 > 1 0.905±0.069 0.146±0.053 < 0 > 1 IRM [3] 0.930±0.012 0.157±0.013 < 0 > 1 < 0 > 1 < 0 > 1 RPGNN [21] 0.929±0.003 0.158±0.004 < 0 > 1 0.914±0.020 0.144±0.017 < 0 > 1 SSR [7] 0.818±0.091 0.248±0.062 < 0 > 1 0.883±0.021 0.168±0.015 < 0 > 1 DISGEN [14] 0.163±0.081 0.542±0.027 0.179±0.082 0.537±0.027 0.232±0.047 0.432±0.013 0.174±0.062 0.447±0.017 BFGCN [22] 0.904±0.014 0.184±0.014 0.883±0.018 0.203±0.016 0.916±0.014 0.137±0.006 0.901±0.014 0.146±0.008 GRIN-RepAug 0.952±0.005 0.129±0.007 0.942±0.006 0.143±0.008 0.946±0.010 0.113±0.013 0.936±0.010 0.125±0.010 GRIN 0.956±0.002 0.125±0.002 0.952±0.001 0.129±0.002 0.952±0.005 0.108±0.006 0.947±0.005 0.113±0.005 GIN GIN [37] 0.918±0.006 0.170±0.006 0.903±0.007 0.185±0.007 0.925±0.004 0.135±0.004 0.914±0.004 0.144±0.003 DIR [35] < 0 > 1 < 0 > 1 < 0 > 1 < 0 > 1 GREA [18] 0.831±0.091 0.237±0.072 < 0 > 1 0.918±0.002 0.141±0.002 < 0 > 1 IRM [3] 0.930±0.010 0.157±0.011 < 0 > 1 0.934±0.003 0.126±0.003 < 0 > 1 RPGNN [21] 0.929±0.001 0.158±0.002 < 0 > 1 0.917±0.025 0.141±0.021 < 0 > 1 SSR [7] 0.810±0.161 0.245±0.104 < 0 > 1 0.860±0.067 0.182±0.042 < 0 > 1 DISGEN [14] 0.123±0.146 0.554±0.048 0.100±0.107 0.565±0.035 0.232±0.047 0.432±0.013 0.181±0.060 0.447±0.016 BFGIN [22] 0.925±0.006 0.163±0.007 0.918±0.004 0.170±0.004 0.929±0.006 0.131±0.006 0.918±0.002 0.141±0.002 GRIN-RepAug 0.954±0.004 0.127±0.006 0.947±0.004 0.137±0.005 0.963±0.001 0.095±0.001 0.953±0.003 0.107±0.004 GRIN 0.961±0.001 0.117±0.002 0.960±0.001 0.119±0.002 0.965±0.001 0.092±0.001 0.962±0.001 0.096±0.001